# Pretraining with random noise for uncertainty calibration

### v1

@@ Abstract
Uncertainty calibration, the process of aligning confidence with accuracy, is a hallmark of human intelligence. However, most machine learning models struggle to achieve this alignment, particularly when the training dataset is small relative to the network’s capacity. Here, we demonstrate that uncertainty calibration can be effectively achieved through a pretraining method inspired by developmental neuroscience. Specifically, training with random noise before data training allows neural networks to calibrate their uncertainty, ensuring that confidence levels are aligned with actual accuracy. We show that randomly initialized, untrained networks tend to exhibit erroneously high confidence, but pretraining with random noise effectively calibrates these networks, bringing their confidence down to chance levels across input spaces. As a result, networks pretrained with random noise exhibit optimal calibration, with confidence closely aligned with accuracy throughout subsequent data training. These pre-calibrated networks also perform better at identifying “unknown data” by exhibiting lower confidence for out-of-distribution samples. Our findings provide a fundamental solution for uncertainty calibration in both in-distribution and out-of-distribution contexts.
One sentence summary
Pretraining with random noise enables networks to align confidence with accuracy, effectively calibrating uncertainty for both in-distribution and out-of-distribution data.
Research Highlights
  • Deep neural networks often struggle to properly calibrate both accuracy and confidence.
  • Pretraining with random noise aligns confidence levels with actual accuracy.
  • Random noise pretraining facilitates meta-learning by aligning confidence with the chance level.
  • Pre-calibrated networks assess both known and unknown datasets through confidence estimation.
@@ Introduction
Fig. 1: Importance of uncertainty calibration in artificailly intelligent systems
(a) The illustration depicts how self-driving cars detect objects in their environment and make decisions based on these detections. A pedestrian wearing a green cap is identified, and both the calibrated and miscalibrated models predict the same label. However, the confidence levels differ significantly between the two models. Miscalibrated confidence can lead to incorrect decisions, even when the predicted label is correct.
@@
Recent advances in deep neural networks have significantly enhanced their ability to learn with accuracy 1. For instance, deep learning models achieved human-level accuracy in classifying natural images over a decade ago 2, 3, 4, 5. Building on this remarkable capability, deep neural networks are now being applied to a wide range of real-world problems that involve complex decision-making, such as autonomous driving systems 6, 7, medical diagnosis 8, 9, and financial engineering 10, 11. However, in these applications, simply inferring the correct answer is insufficient; an additional dimension of information must be considered: confidence or uncertainty (Fig. 1). It is crucial to predict answers with calibrated probabilities that accurately reflect the likelihood of correctness.
Despite their remarkable capacity for accurate learning, modern deep neural networks are often poorly calibrated in terms of confidence estimation 12, 13, 14. Specifically, the rapid increase in model capacity enables accurate classification but negatively impacts confidence calibration 12. Modern neural networks frequently exhibit overconfidence, even with out-of-distribution samples despite lacking knowledge about them 15, 16. Miscalibrated confidence poses significant problems in real-world applications, where false predictions can be critical and costly 17, 18, 19, 20, 21, 22. Hallucinations in large language models 23, 24, 25, 26, where the model confidently produces false and unsubstantiated outputs 27, 28, 29, 30, are a well-known issue arising from the miscalibration of neural networks 31, 32, 33.
Several studies have proposed solutions to calibrate confidence for predictions from classifier models. These research efforts suggest post-processing techniques to compute calibrated probabilities from the naïve output of the trained network, such as binning the output probability 34, 35, 36, 12 or adjusting the level of network output 36, 37, 12. Recent studies also utilize Bayesian neural networks to achieve state-of-the-art performance in uncertainty calibration 38, 39, 40, but they require computationally intensive repetitions to calculate uncertainty in each trial. Many state-of-the-art algorithms employ auxiliary networks or score functions to distinguish overconfident predictions for unseen out-of-distribution data from those for in-distribution data 15, 16, 41, 42, 43, 44. Overall, previous research typically handles calibration issues separately for in-distribution data and unseen out-of-distribution data. Moreover, these methods often require additional, computationally expensive steps to obtain calibrated confidence or auxiliary scores.
@@
In contrast to the scenario in machine learning, where calibrated probabilities must be obtained through additional processes, human intelligence naturally exhibits meta-cognition, which estimates uncertainty and is aware of what is “known” and “unknown” 45, 46, 47, 48. This raises the question: what is the underlying mechanism of uncertainty calibration in the biological brain, and how can we enable artificial neural networks to learn human-like uncertainty calibration without relying on post-processing or auxiliary computation? To address this question, we draw inspiration from our previous work 49, which proposed pretraining with random noise as a method that mirrors the biological prenatal learning stage. Specifically, we focused on the fact that random noise pretraining stabilizes the input space to align with the output chance level uniformly before learning from data. We hypothesized that these initial calibrations could facilitate the learning of calibrated uncertainty during subsequent data training.
@@
In this study, we demonstrate that pre-calibrating uncertainty through training with random noise enables neural networks to learn probabilities that accurately reflect the likelihood of correctness. Our findings show that randomly initialized neural networks tend to exhibit high confidence even when they lack sufficient knowledge, but training with random noise effectively reduces loss by calibrating this confidence to chance levels uniformly across input spaces. As a result, pre-calibrated networks learn calibrated probabilities with accuracy and confidence aligned throughout subsequent data training. Moreover, pre-calibrated networks exhibit lower confidence when encountering unknown data, aiding in the detection of out-of-distribution samples. Taken together, our results suggest pretraining with random noise as a simple yet powerful strategy to address uncertainty calibration issues.
@@ Results
Fig. 2: Neural network complexity and data size as co-determinants for calibration
(a) The predicted answer and its confidence are calculated using the probability output from the SoftMax function. (b) Reliability diagram for a 2-layer feedforward neural network trained on a subset of the CIFAR-10 dataset (training data size = 1000). The test dataset’s predictive confidence and correctness are binned based on confidence values, and the accuracy is calculated in each bin. The diagonal line represents ideal calibration, where the confidence perfectly matches the expected accuracy. The Expected Calibration Error (ECE) is the difference between predicted confidence and actual accuracy. (c) Calibration error across various model complexities and training data sizes. Each color represents the ECE: blue and red indicate low and high calibration error, respectively. (d) The effect of reducing the training data size on calibration error, with model complexity fixed (depth = 2). (e) The effect of increasing model complexity on calibration error, with training data size fixed (training data size = 1000).
@@
To investigate the issue of confidence calibration in deep neural network models, we used a feedforward neural network designed for pattern classification of three-channel natural images (32×32×3) into ten categories. The network consists of multiple hidden layers with ReLU nonlinearities and a final classification layer that employs the SoftMax function. For a given image sample, the model outputs probabilities for each object category (Fig. 2a), and the predicted label corresponds to the category with the maximum probability. The model also provides a measure of prediction confidence, which is derived from the probability assigned to the predicted category. The predictive uncertainty, can be quantified as the difference between the confidence value and one. Ideally, a neural network is expected to output calibrated confidence, where the confidence level accurately reflects the actual likelihood of the prediction being correct. In contrast, a poorly calibrated network may output confidence levels that do not correspond to the actual likelihood of correctness, leading to potential misinterpretations of the model’s predictions.
@@
We trained the feedforward neural network using a subset of the CIFAR-10 dataset 50, which consists of natural images labeled with ten classes representing animals and objects. After training, we evaluated the model’s predictions and confidence using test images that were not part of the training set. Specifically, we investigated whether the confidence level of the trained network accurately reflects the likelihood of correctness. By measuring binned confidence values and the corresponding expected accuracy within each binned trial prediction, we constructed a reliability diagram 51, 37 to visualize the model’s calibration (Fig. 2b). In this diagram, ideal calibration appears as an identity function, where the estimated confidence corresponds to the actual probability of correctness. In our results, we observed a significant gap between ideal calibration and the model’s outputs. In general, the model’s accuracy was lower than its confidence level, demonstrating the model’s tendency for overconfidence.
@@
We investigated the pattern of miscalibration by training neural networks of varying complexity with different training data sizes. To quantitatively measure the degree of miscalibration, we calculated the expected calibration error (ECE) 52, which averages the difference between accuracy and confidence in each bin, weighted by the number of samples. As a result, we found that both model complexity and training data size systematically affect confidence calibration (Fig. 2c). Specifically, we observed that the degree of miscalibration increases as the training data size decreases (Fig. 2d) and as model complexity increases (Fig. 2e). Thus, miscalibration arises when the training data is insufficient relative to the network’s complexity. This result explains why state-of-the-art models with high complexity struggle with the mismatch between confidence and accuracy. Recent network architectures often increase in structural depth to enhance learning capacity, but the available data size does not increase correspondingly. In real-world applications, acquiring and training on massive datasets is expensive and requires substantial computational resources. As a result, the miscalibration problem becomes inevitable in deep learning models.
@@
Fig. 3: Pretraining with random noise enables confidence calibration of neural networks
@@
(a) Schematic illustrating the pretraining with random noise 49 and subsequent training with real data. During the random noise pretraining stage, the network is trained on randomly sampled inputs from a Gaussian distribution and randomly sampled labels from a uniform distribution, which are not paired. (b) Test loss across random noise pretraining and data training. The blue curve represents networks pretrained with random noise and subsequently trained with real data, while the orange curve represents networks trained solely with data. (c) Test accuracy during data training. (d) Histogram of confidence for a network trained only with data (orange) and a network pretrained with random noise (blue). Averaged accuracy and confidence are indicated by vertical lines. (e) The difference between averaged confidence and accuracy of predictions. (f) Reliability diagram showing the expected accuracy for samples binned by confidence. (g) Expected calibration error. (h) Effect of random noise pretraining on calibration error as training data size is reduced, with model complexity fixed (depth = 2). (i) Effect of random noise pretraining on calibration error as model complexity increases, with training data size fixed (training data size = 1000). (j) The combined effect of random noise pretraining under varying conditions of training data size and network complexity on calibration error.
@@
Recently, we introduced a novel strategy for pretraining networks inspired by the development of the biological brain 49. In this work, we demonstrated that pretraining networks with random noise enables robust learning without weight transport. Here, we revisit our random noise pretraining method and show that this process facilitates the preconditioning of networks, calibrating confidence to the chance level. We pretrained neural networks with noise inputs randomly sampled from a Gaussian distribution, and with random labels sampled from a uniform distribution (Fig. 3a, Supplementary Fig. 1). Notably, we observed that the loss gradually decreased when the network was trained with random noise, even though the accuracy remained at the chance level (Fig. 3b, Supplementary Fig. 2). Furthermore, when the network was trained with real data following the noise pretraining (Fig. 3b, Supplementary Fig. 3), we observed a more robust decrease in test loss during subsequent training, which was significantly lower than in networks trained without random noise pretraining (Fig. 3b). We also confirmed that the network pretrained with random noise achieved higher accuracy (Fig. 3c). These results demonstrate that pretraining with random noise enables faster and more effective loss reduction in subsequent training.
@@
Next, we examined whether random noise pretraining can improve the calibration of probabilities to align confidence with accuracy levels. First, we measured the confidence distribution of the network’s output on the test dataset and compared it with the actual accuracy (Fig. 3d). We found that random noise pretraining significantly reduced the difference between confidence and accuracy in subsequent data training (Fig. 3e, w/o vs. w/ random pretraining, $n_{\text{net}}=10$, Wilcoxon rank-sum test, $P<10^{-3}$; w/o vs. zero, Wilcoxon signed-rank test, NS, $P<10^{-3}$; w/ vs. zero, Wilcoxon signed-rank test, NS, $P=0.131$). We also investigated the difference between the two networks using a reliability diagram (Fig. 3f). We observed that the randomly pretrained network’s outputs were closer to ideal calibration than those of networks trained with data alone. Additionally, we confirmed that the pretrained networks exhibited significantly lower calibration error compared to networks solely trained with data (Fig. 3g, w/o vs. w/ random pretraining, $n_{\text{net}}=10$, Wilcoxon rank-sum test, $P<10^{-3}$).
We then extended this analysis to neural networks trained with varying amounts of data and found that the effect of random noise pretraining in reducing calibration error was significant, regardless of the training data size (Fig. 3h, w/o vs. w/ random pretraining, $n_{\text{net}}=10$, Wilcoxon rank-sum test, $P<10^{-3}$). We also tested neural networks of different depths and confirmed the robust effect of noise pretraining (Fig. 3h, w/o vs. w/ random pretraining, $n_{\text{net}}=10$, Wilcoxon rank-sum test, $P<10^{-3}$). Furthermore, we measured calibration error for various combinations of training data size and network depth (Fig. 3j) and found that random noise pretraining successfully calibrated the confidence level across the conditions we examined. The effect of pretraining was particularly significant for networks with deep structures trained on smaller datasets, where confidence miscalibration typically occurs.
In addition to the current results using the standard training method with backpropagation, we also investigated neural networks trained with a biologically plausible algorithm, employing feedback alignment (Supplementary Fig. 4). We found that miscalibration also occurs in this type of model, suggesting that the issue of miscalibration is not limited to the backpropagation algorithm but may be a universal challenge across various algorithms. Notably, we observed that random noise pretraining significantly improves confidence calibration, even in this case, indicating that random noise pretraining could provide a general solution for confidence calibration across different models.
@@
Pre-calibration of network uncertainty over input space via random noise
@@
Fig. 4: Random noise pre-calibrates neural network uncertainty over input space
@@
(a) Diagram illustrating the training of random noise in a toy-model network with a two-dimensional input space and binary output space. The network receives random noise with two input features and randomly assigned binary labels, and is trained using Binary Cross-Entropy (BCE) loss. (b) Visualization of confidence over the input space. (Left) Confidence map of the untrained network. (Right) Confidence map after random noise pretraining. (c) Confidence distribution of the network for two-dimensional noise. (d) Class bias of the network for two-dimensional noise, showing the extent of bias in class prediction. (e) Diagram of training random noise in a multi-layer perceptron with a high-dimensional input space, designed to classify the CIFAR-10 dataset. The network receives 32×32×3 input images and outputs SoftMax probabilities for ten classes. (f) Visualization of SoftMax output probabilities for random noise inputs in the readout neurons. (Left) Untrained network. (Right) Network after random noise pretraining. The vertical line represents the chance level of classification. (g) Confidence distribution of the network for high-dimensional noise. (h) Class bias of the network for high-dimensional noise, showing the bias towards specific classes. (i-j) Confidence reduction for various unseen real data during random noise pretraining. (i) CIFAR-10 and SVHN datasets presented to the network during random noise training, with measured accuracy and confidence (Left). Confidence and accuracy for the CIFAR-10 dataset (Right). (j) Confidence distribution for CIFAR-10 and SVHN datasets before and after random noise pretraining. (k) Expected Calibration Error (ECE) for CIFAR-10 and SVHN datasets before and after random noise pretraining.
We further investigated how pretraining with random noise influences the properties of neural networks prior to their exposure to real data. To explore this, we introduced a simplified toy-model network with a two-dimensional input space (Fig. 4a), enabling straightforward visualization of the input space with only two features (Fig. 4b). Under these conditions, we observed that the initial state of a conventionally processed network — randomly initialized 53 and untrained — exhibited a highly biased and arbitrary distribution of confidence across the input space (Fig. 4b, Left). Additionally, we confirmed that this initial miscalibration was consistent across multiple trials with randomly initialized networks (Supplementary Fig. 5a).
@@
In contrast, networks pretrained with random noise exhibited a more homogeneous distribution of confidence across the input space, effectively calibrated to the chance level (Fig. 4b, Right; Supplementary Fig. 5b). This result demonstrates that pretraining with random noise mitigates overconfidence by calibrating the confidence levels across the entire input space, bringing them closer to the chance level (Fig. 4c, w/ vs. w/o random pretraining, $n_{\text{trial}}=1000$, Wilcoxon rank-sum test, $P<10^{-3}$; Supplementary Fig. 5c). We also observed that untrained networks could exhibit initial biases toward specific output classes due to fluctuations in confidence. However, random noise pretraining significantly reduced this initial class bias, resulting in a more uniform confidence distribution across the classes (Fig. 4d, w/ vs. w/o random pretraining, $n_{\text{net}}=10$, Wilcoxon rank-sum test, $P<10^{-3}$).
@@
Next, we extended this analysis to models with higher-dimensional input spaces (Fig. 4e). Specifically, the model used input features with dimensions of 32×32×3 and output Softmax probabilities for ten classes, allowing it to train on CIFAR-10 data. We first measured the probability outputs for random noise inputs in both untrained and randomly pretrained networks (Fig. 4f). Consistent with the toy model results, we observed overconfidence and class bias in untrained networks (Fig. 4f, Left, $n_{\text{trial}}=100$, one-way ANOVA, $P<10^{-3}$), whereas such biases were barely detectable in randomly pretrained networks (Fig. 4f, Right, $n_{\text{trial}}=100$, one-way ANOVA, NS, $P=0.429$). We confirmed that the differences between the two networks were consistent with the toy model findings, both in the distribution of confidence (Fig. 4g, w/ vs. w/o random pretraining, $n_{\text{trial}}=100$, Wilcoxon rank-sum test, $P<10^{-3}$) and in class bias (Fig. 4h, w/ vs. w/o random pretraining, $n_{\text{net}}=10$, Wilcoxon rank-sum test, $P<10^{-3}$). These results suggest that random noise pretraining enables the network to maintain lower, more uniformly calibrated confidence levels, even when handling more complex, higher-dimensional input data.

### v2

@@ Abstract
Uncertainty calibration is crucial for various machine learning applications, yet it remains challenging. Many models exhibit hallucinations — confident yet inaccurate responses — due to miscalibrated confidence. Here, we show that the common practice of random initialization in deep learning, often considered a standard technique, is an underlying cause of this miscalibration, leading to excessively high confidence in untrained networks. Our method, inspired by developmental neuroscience, addresses this issue by simply pretraining networks with random noise and labels, reducing overconfidence and bringing initial confidence levels closer to chance. This ensures optimal calibration, aligning confidence with accuracy during subsequent data training, without the need for additional pre- or post-processing. Pre-calibrated networks excel at identifying “unknown data,” showing low confidence for out-of-distribution inputs, thereby resolving confidence miscalibration.
@@ Introduction
Recent advances in deep neural networks have significantly improved their ability to learn with high accuracy 1. For example, deep learning models achieved human-level accuracy in classifying natural images over a decade ago 2, 3, 4, 5. These models are now being applied to a wide range of real-world problems that involve complex decision-making, such as autonomous driving 6, medical diagnosis 7, and financial engineering 8. In these applications, simply inferring the correct answer is insufficient; an additional dimension of information, confidence or uncertainty, is crucial to predict answers with calibrated probabilities that accurately reflect the likelihood of correctness (Fig. 1a). However, modern deep neural networks are often poorly calibrated for confidence estimation 9, 10, 11. Ironically, the increase in model capacity enables accurate performance but often negatively impacts confidence calibration 9. Specifically, many recent models frequently exhibit overconfidence, even with out-of-distribution samples, despite lacking knowledge about them 12, 13. This miscalibration leads to poor decision-making in real-world applications 14, 15 — For example, hallucinations in large language models 16, 17, where the model confidently produces false and unsubstantiated outputs 18, 19, are a well-known issue resulting from the miscalibration of neural networks 20, 21.
Fig. 1: Confidence miscalibration in artificial neural networks.
(a) The illustration depicts how self-driving cars detect objects in their environment and make decisions based on these detections. Both the calibrated and miscalibrated models predict the same label, but the confidence levels differ significantly between the two models. Miscalibrated confidence can lead to incorrect decisions, even when the model’s accuracy is high. (b) The predicted answer and its confidence are calculated using the probability output from the SoftMax function. (c) Reliability diagram for a six-layer feedforward neural network trained on a subset of the CIFAR-10 dataset (training data size = 4,000). The test dataset’s predictive confidence and correctness are binned based on confidence values, and the accuracy is calculated in each bin. The diagonal line represents ideal calibration, where confidence perfectly matches the expected accuracy. The Expected Calibration Error (ECE) is the difference between predicted confidence and actual accuracy. (d) Calibration error across various model complexities and training data sizes. Each color represents the ECE: blue and red indicate low and high calibration error, respectively.
Previous studies have tried to address calibration issues through additional pre- or post-processing techniques. However, these methods require additional, computationally expensive steps to obtain calibrated confidence 9, 22, 23, 24, 25. Many state-of-the-art algorithms employ auxiliary networks or score functions to differentiate between overconfident predictions for unseen out-of-distribution data and those for in-distribution data 26, 27, 28. In general, the fundamental problem is that previous approaches typically treat in-distribution and out-of-distribution data separately, which essentially recalibrates the confidence values through separate calculation processes to achieve the desired results.
@@
We can ask a more fundamental question: What causes uncertainty miscalibration in deep neural networks? Unlike machines, human intelligence naturally exhibits metacognition, enabling the estimation of uncertainty and awareness of what is “known” and “unknown” 29, 30, 31, 32, 33. So, what is the underlying mechanism of uncertainty calibration in the brain, and how can artificial neural networks learn brain-like uncertainty calibration, possibly without post-processing or auxiliary computation? To address this, we draw inspiration from our previous work 34, which proposed pretraining with random noise as a method that mimics the biological prenatal learning stage. Specifically, we hypothesize that these initial pre-calibrations could facilitate the learning of calibrated uncertainty.
@@
In this study, we demonstrate that the commonly used method of random initialization in deep learning, long regarded as a standard technique, is, in fact, the underlying cause of network uncertainty miscalibration. The solution to this miscalibration issue is achieved simply through a novel initialization strategy that deviates from conventional practices — by pretraining networks using random noise and labels, without additional pre- or post-processing methods. Our findings show that randomly initialized neural networks often exhibit high confidence even when they lack sufficient knowledge, but training with random noise effectively calibrates this confidence to uniform chance levels across the input spaces. As a result, the networks learn calibrated probabilities, with accuracy and confidence aligned throughout subsequent data training. Additionally, pre-calibrated networks show lower confidence when encountering unknown data, which helps in detecting out-of-distribution samples. Overall, our results suggest that pretraining with random noise is a simple yet powerful strategy for addressing uncertainty calibration issues.
@@ Results
To investigate the pattern of confidence miscalibration in deep neural network models, we first used a feedforward neural network designed for pattern classification of three-channel natural images (32×32×3) into ten categories (Fig. 1b) (see Methods for details). The network consists of multiple hidden layers with ReLU nonlinearities and a final classification layer that employs the SoftMax function. For a given image sample, the model outputs probabilities for each object category, and the predicted label corresponds to the category with the maximum probability. The model also provides a measure of prediction confidence, which is derived from the probability assigned to the predicted category. The predictive uncertainty can be quantified as the difference between the confidence value and one. Ideally, a neural network is expected to output calibrated confidence, where the confidence level accurately reflects the actual likelihood of the prediction being correct (Fig. 1c and Supplementary Fig. 1). In contrast, a poorly calibrated network may output confidence levels that do not correspond to the actual likelihood of correctness, leading to potential misinterpretations of the model’s predictions.
@@
We trained the feedforward neural network using a subset of the CIFAR-10 dataset 35, which consists of natural images labeled with ten classes representing animals and objects. After training, we evaluated the model’s predictions and confidence using test images that were not part of the training set. Specifically, we investigated whether the confidence level of the trained network accurately reflects the likelihood of correctness (Supplementary Fig. 1). By measuring binned confidence values and the corresponding expected accuracy within each binned trial prediction, we constructed a reliability diagram 36, 22 to visualize the model’s calibration (Fig. 1c). In this diagram, ideal calibration appears as an identity function, where the estimated confidence corresponds to the actual probability of correctness.
@@
We investigated the pattern of miscalibration by training neural networks of varying complexity with different training data sizes (Fig. 1d). To quantitatively measure the degree of miscalibration, we calculated the Expected Calibration Error (ECE) 37, which averages the difference between accuracy and confidence in each bin, weighted by the number of samples. In general, we observed a significant gap between ideal calibration and the model’s outputs — the model’s accuracy was lower than its confidence level, demonstrating the model’s tendency for overconfidence in most conditions. We found that both model complexity and training data size systematically affect confidence calibration. Specifically, the degree of miscalibration increases as the training data size decreases and as model complexity increases, suggesting that miscalibration occurs when the training data is insufficient relative to the network’s complexity. This result explains why state-of-the-art models with high complexity struggle with the mismatch between confidence and accuracy. Recent network architectures often increase in depth to enhance learning capacity, but the available data size does not scale accordingly. In real-world applications, acquiring large datasets is costly and computationally demanding, making miscalibration inevitable in deep learning models.
@@
Fig. 2: Pretraining with random noise enables confidence calibration in neural networks.
@@
(a) Prenatal learning through spontaneous neural activity prior to sensory input in a fetal rat, before birth (left) and after eye opening (right) 38. (b) Spontaneous neural activity in the developing visual and auditory areas 39. (c) Schematic of the pretraining algorithm with random noise, inspired by the developing brain before sensory experience. During pretraining, the network is trained on randomly sampled inputs from a Gaussian distribution and unpaired labels from a uniform distribution. (d) Test loss during random noise pretraining and data training. (Inset) Accuracy during data training in networks with (blue) and without (orange) pretraining. (e) Histogram of confidence for a network trained only with data (orange) and a network pretrained with random noise (blue). Averaged accuracy and confidence are indicated by vertical lines. (Right) The difference between averaged confidence and accuracy of predictions. (f) Reliability diagram showing the expected accuracy for samples binned by confidence. (Inset) Expected Calibration Error. (g) The effect of random noise pretraining under varying conditions of training data size and network complexity on calibration error. The marker indicates the parameters used in (d-f), where the six-layer networks are trained with a dataset size of 4,000.
The biological brain begins its learning process even before birth, prior to encountering sensory inputs, through spontaneous neural activity 38 (Fig. 2a). During this early stage, the brain optimizes its neural connections by transmitting this spontaneous activity to higher cortical regions, and this prenatal learning is observed across multiple brain areas involved in sensory systems 40, 41, 38, 39 (Fig. 2b). In particular, the spontaneous activities take on various spatiotemporal forms, such as waves 40 or local synchronization 41. However, when considered as a whole, they exhibit statistically random patterns. Thus, this process can be seen as a form of pretraining with random noise 34. Here, we demonstrate that a pretraining algorithm inspired by this biological process facilitates the preconditioning of networks, calibrating their confidence to a chance level.
@@
We pretrained neural networks using noise inputs randomly sampled from a Gaussian distribution and random labels sampled from a uniform distribution (Fig. 2c). During training with random noise, we observed that the loss gradually decreased, while accuracy remained at the chance level (Fig. 2d; Supplementary Fig. 2). When the network was subsequently trained with real data after the noise pretraining, we observed a more substantial reduction in test loss, which was significantly lower compared to networks trained without random noise pretraining. Additionally, the network pretrained with random noise achieved higher accuracy (Fig. 2d, Inset; Supplementary Fig. 2). These results demonstrate that pretraining with random noise facilitates faster and more effective loss reduction during subsequent training with real data.
@@
Next, we investigated whether random noise pretraining could improve the calibration of probabilities, aligning confidence with accuracy. We began by measuring the confidence distribution of the network’s output on the test dataset and compared it to the actual accuracy (Fig. 2e). We found that random noise pretraining significantly reduced the difference between confidence and accuracy during subsequent data training (Fig. 2e, Right, w/o vs. w/ random pretraining, $n_{\text{net}}=10$, Wilcoxon rank-sum test, $P<10^{-3}$; w/o vs. zero, Wilcoxon signed-rank test, NS, $P<10^{-3}$; w/ vs. zero, Wilcoxon signed-rank test, NS, $P=0.432$). We further examined this difference using a reliability diagram (Fig. 2f), which showed that the network pretrained with random noise had outputs that were closer to ideal calibration than networks trained on data alone. Additionally, we confirmed that the pretrained networks exhibited significantly lower calibration error compared to those trained only with data (Fig. 2f, Inset, w/o vs. w/ random pretraining, $n_{\text{net}}=10$, Wilcoxon rank-sum test, $P<10^{-3}$). We then extended this analysis to networks with varying depths and sizes of training data. We found that random noise pretraining had a significant effect, regardless of network depth or training data size (Fig. 2g, w/o vs. w/ random pretraining, $n_{\text{net}}=10$, Wilcoxon rank-sum test, $P<10^{-3}$). Pretraining successfully reduced calibration error across all conditions examined. Notably, the effect was particularly pronounced for networks with deeper structures trained on smaller datasets, where confidence miscalibration is more common. In addition to the results obtained using the standard backpropagation 42 training method, we also examined networks trained with a biologically plausible algorithm employing feedback alignment 43 (Supplementary Fig. 3). We observed that miscalibration also occurred in this model, but random noise pretraining significantly improved confidence calibration. This finding suggests that random noise pretraining could serve as a general solution for enhancing confidence calibration across various network learning models.
@@
Pre-calibration of network uncertainty across the input space using random noise
@@
Fig. 3: Random noise pre-calibrates neural network uncertainty over input space.
@@
(a) Diagram illustrating the training of random noise in a toy-model network with a two-dimensional input space and binary output space. The network receives random noise with two input features and randomly assigned binary labels, and is trained using Binary Cross-Entropy (BCE) loss. (b) Visualization of confidence over the input space. (Left) Confidence map of the untrained network. (Right) Confidence map after random noise pretraining. (c) Confidence distribution of the network for two-dimensional noise. (d) Class bias of the network for two-dimensional noise, showing the extent of bias in class prediction. (e) Diagram of training random noise in a multi-layer perceptron with a high-dimensional input space, designed to classify the CIFAR-10 dataset. The network receives 32×32×3 input images and outputs SoftMax probabilities for ten classes. (f) Visualization of SoftMax output probabilities for random noise inputs in the readout neurons. (Left) Untrained network. (Right) Network after random noise pretraining. The dashed line represents the chance level of classification. (g) Confidence distribution of the network. (h) Class bias of the network, showing the bias towards specific classes.
@@
To better understand how pretraining with random noise affects confidence calibration, we further explored the characteristics of randomly initialized, untrained networks and those that were randomly pretrained, prior to exposure to real data. For this purpose, we introduced a simplified toy-model network with a two-dimensional input space (Fig. 3a), allowing for visualization of the input feature space (Fig. 3b). Under these conditions, we observed that the initial state of a conventionally processed network — randomly initialized 44 and untrained — displayed a highly biased distribution of confidence across the input space (Fig. 3b, Left). Moreover, we confirmed that this initial miscalibration was consistent across multiple trials with randomly initialized networks (Supplementary Fig. 4a). In contrast, networks pretrained with random noise displayed a more homogeneous confidence distribution across the input space, effectively calibrating to the chance level (Fig. 3b, Right; Supplementary Fig. 4b). This result demonstrates that pretraining with random noise helps mitigate overconfidence by calibrating confidence levels across the entire input space, bringing them closer to chance (Fig. 3c, w/ vs. w/o random pretraining, $n_{\text{trial}}=1000$, Wilcoxon rank-sum test, $P<10^{-3}$; Supplementary Fig. 4c). We also observed that untrained networks could exhibit initial biases toward specific output classes due to fluctuations in confidence. However, random noise pretraining significantly reduced this initial class bias, leading to a more uniform confidence distribution across the classes (Fig. 3d, w/ vs. w/o random pretraining, $n_{\text{net}}=10$, Wilcoxon rank-sum test, $P<10^{-3}$).
@@
Next, we extended this analysis to models with higher-dimensional input spaces (Fig. 3e). Specifically, we used a model with input features of 32×32×3 dimensions and output SoftMax probabilities for ten classes, allowing it to train on CIFAR-10 data. We first measured the probability outputs for random noise inputs in both untrained and randomly pretrained networks (Fig. 3f; Supplementary Fig. 5). Consistent with the toy model results, we observed overconfidence and class bias in untrained networks (Fig. 3f, Left, $n_{\text{trial}}=100$, one-way ANOVA, $P<10^{-3}$; Supplementary Fig. 5a), whereas such biases were barely detectable in randomly pretrained networks (Fig. 3f, Right, $n_{\text{trial}}=100$, one-way ANOVA, NS, $P=0.429$; Supplementary Fig. 5b). We confirmed that these differences aligned with the toy model results, both in terms of confidence distribution (Fig. 3g, w/ vs. w/o random pretraining, $n_{\text{trial}}=100$, Wilcoxon rank-sum test, $P<10^{-3}$; Supplementary Fig. 5c) and class bias (Fig. 3h, w/ vs. w/o random pretraining, $n_{\text{net}}=10$, Wilcoxon rank-sum test, $P<10^{-3}$). These results suggest that random noise pretraining enables the network to maintain lower, more uniformly calibrated confidence levels, even when handling more complex, higher-dimensional input data.